# First principles band structure of interacting phosphorus and boron/aluminum $\delta$-doped layers in silicon

### v1

@@ III.4 Considerations when comparing DFT results to experimental systems
It should be noted that the DFT simulations provided in this work are necessarily an approximation of realistic systems with both acceptor and donor $\delta$-doped layers. Many of the errors are unfortunately inherent to any DFT-based approach for materials simulation of electronic properties. The band gap of the systems will be somewhat underestimated, an effect that has been well documented in various DFT calculations. We apply SCAN pseudopotentialsSun et al. (2015); Yao and Kanai (2017) to combat this problem, but it has not been entirely solved, as evidenced by our prediction of a pure silicon supercell of 0.98 eV. This may be particularly relevant at separation distances leading to us predicting a band structure that is continuous across the typical silicon band gap, such as the B-P 100 Å case (Fig. 3c). There may, in fact, be slight band gaps in these materials that we cannot see without significantly more computationally expensive methods. Furthermore, there is some uncertainty in the exact placement of the Fermi level, which would be dependent on the specific doping of any given sample as well as the measurement temperature. When the Fermi level is placed firmly within the typical silicon band gap (as in the 1 and 4 Å separation cases for both B-P and Al-P layers), it is particularly ill-defined within DFT and should be taken more clearly as an indication that the valence band is fully occupied and the conduction band is fully unoccupied.
There are three additional aspects of the results where the approximations necessary for DFT calculations become especially notable. The first is in the concentration effects of using a limited DFT supercell with periodic boundary conditions, which necessarily forces the concentration of dopants being modeled to be relatively high. While this high concentration is laterally correct, i.e. roughly 1/4 monolayer coverage, it is unlikely to be correct perpendicular to the $\delta$ layer. We are in fact modeling a silicon system with periodic $\delta$ layers every 218 Å  rather than a single isolated pair of $\delta$ layers. As recently explored by Campbell et al., Campbell et al. (2023) it can be difficult to fully disentangle the impact of $\delta$-doped layers from the high concentration of dopants being simulated. Nonetheless, given the agreement between experimental characterization and DFT calculations of phosphorus $\delta$-doped layers,Miwa et al. (2013); Mazzola et al. (2014a, b); Miwa et al. (2014); Mazzola et al. (2018); Holt et al. (2020) it is reasonable to take these DFT results as good first-order approximations of the electronic structure.
@@ IV Conclusion
Motivated by recent progress in using APAM to create $\delta$-doped layers, we used DFT to predict the electronic structure of both interacting phosphorus and boron $\delta$-doped layers and interacting phosphorus and aluminum $\delta$-doped layers in silicon. At separations of less than $\approx$ 10 Å, we demonstrate that these $\delta$-doped layers will strongly suppress each other, leading to an effective band gap and electronic structure close to that of bulk silicon. As the separation between the layers increases to around 10 Å, the suppression decreases, and distinct $\delta$ potentials can be seen at the site of each layer, essentially creating a p-n diode with an intrinsic layer of silicon in the place of the depletion region. This decrease in suppression manifests as a change in the effective band gap that can be tailored by controlling the amount of silicon between the $\delta$-doped layers, ranging from 0.36 eV for 20 Å spacing to 0.86 eV for 1 Å spacing. However, the CBM and VBM are spatially separated, so carrier transport between $\delta$-doped layers likely requires tunneling or application of bias. Calculations of the $\delta$-doped layer potentials enable us to predict the tunneling rate between the two layers, finding that the tunneling at higher separations is greater than what would be expected for a pure 1.1 eV triangular barrier. These calculations thus provide a foundation for design of silicon electronics based on interacting $\delta$ layers.
@@ Appendix A Band Structures of isolated single $\delta$-doped layers
The band structure and LDOS of a phosphorus $\delta$-doped layer is shown in Fig. 8, a boron $\delta$-doped layer band structure is shown in Fig. 9, and an aluminum $\delta$-doped layer band structure is shown in Fig. 10. For each structure, the effects are limited entirely to the conduction band (in the case of phosphorus) or the valence band (in the case of boron). Naively, we might expect that the effects of including both a phosphorus and boron $\delta$-doped layer would be a superposition of the phosphorus conduction band and the boron conduction bands. This description only resembles the band structures when the $\delta$-doped layers are separated by $\geq$ 20 Å, indicating that interaction between the layers is indeed causing significant suppression of the bands at lower separations. It is notable that SCAN pseudopotentials predict that the $\delta$ potential induced by the phosphorus $\delta$-doped layer is much more significant in magnitude than previous predictions. While the exact magnitude is likely an overestimate, it does indicate that phosphorus $\delta$-doped layers should have an extremely large profile within our calculations. This indicates that the suppression of the phosphorus $\delta$-potential at low separation distances is even more significant than it may appear at first glance.
@@
Figure 11: The doping potential of structures with (a) 1 Å, (b) 4 Å, (c) 10 Å separation between the boron and phosphorus $\delta$-doped layers. The left column displays the potential before the atomic structures are allowed to relax to respond to stress, and the right column is the potential after the structures are allowed to relax.
Figure 12: The doping potential of structures with (a) 20 Å, (b) 100 Å separation between the boron and phosphorus $\delta$-doped layers. The left column displays the potential before the atomic structures are allowed to relax to respond to stress, and the right column is the potential after the structures are allowed to relax.
To further elucidate the nature of the $\delta$-doped layers interaction, we plot the doping potentials for each separation distance in Figs. 11 and  12. These potentials are calculated by subtracting the planar averaged electrostatic potential of a pure silicon supercell from the planar averaged electrostatic potential of the system with boron and phosphorus $\delta$-doped layers. The doping potentials calculated in the left column of each figure represent the potential before the system is allowed to relax in response to the stress that the doping atoms induce in the system. These graphs more or less follow the pattern that would be expected from the LDOS calculated in Figs. 2 and 3. In the right column, however, we display the potential after the system is allowed to relax in response to stress induced by the doping atoms. These potentials have a much more oscillatory character around the $\delta$-doped layers, which is a direct result of atomic movement in response to the strain slightly altering the periodicity of the nearby silicon atoms. These oscillations typically extend $\sim$ 15 Å on either side of the $\delta$-doped layers, indicating the range at which $\delta$-doped layer induced stress can be expected to affect nearby silicon atoms. Similar to what is seen in Ref. Campbell et al., 2023, boron $\delta$-doped layers feature particularly strong oscillation with the resulting $\delta$ potential being less well defined within the material compared to the phosphorus $\delta$-doped layers. This is especially clear in the well separated structures at 20 and 100 Å separation shown in Fig. 12.
Figure 13: The doping potential of Al-P structures with (a) 1 Å, (b) 4 Å, and (c) 10 Å separation between the boron and phosphorus $\delta$-doped layers. The left column displays the potential before the atomic structures are allowed to relax to respond to stress, and the right column is the potential after the structures are allowed to relax.
Figure 14: The doping potential of Al-P structures with (a) 20 Å, (b) 100 Å separation between the boron and phosphorus $\delta$-doped layers. The left column displays the potential before the atomic structures are allowed to relax to respond to stress, and the right column is the potential after the structures are allowed to relax.

### v2

@@ III.4 Considerations when comparing DFT results to experimental systems
It should be noted that the DFT simulations provided in this work are necessarily an approximation of realistic systems with both acceptor and donor $\delta$-doped layers. Many of the errors are unfortunately inherent to any DFT-based approach for materials simulation of electronic properties. The band gap of the systems will be somewhat underestimated, an effect that has been well documented in various DFT calculations. We apply SCAN pseudopotentialsSun et al. (2015); Yao and Kanai (2017) to combat this problem, but it has not been entirely solved, as evidenced by our prediction of a pure silicon supercell of 0.98 eV. This may be particularly relevant at separation distances leading to us predicting a band structure that is continuous across the typical silicon band gap, such as the B-P 10 nm case (Fig. 3c). There may, in fact, be slight band gaps in these materials that we cannot see without significantly more computationally expensive methods. Furthermore, there is some uncertainty in the exact placement of the Fermi level, which would be dependent on the specific doping of any given sample as well as the measurement temperature. When the Fermi level is placed firmly within the typical silicon band gap (as in the 0.1 and 0.4 nm separation cases for both B-P and Al-P layers), it is particularly ill-defined within DFT and should be taken more clearly as an indication that the valence band is fully occupied and the conduction band is fully unoccupied.
There are three additional aspects of the results where the approximations necessary for DFT calculations become especially notable. The first is in the concentration effects of using a limited DFT supercell with periodic boundary conditions, which necessarily forces the concentration of dopants being modeled to be relatively high. While this high concentration is laterally correct, i.e. roughly 1/4 monolayer coverage, it is unlikely to be correct perpendicular to the $\delta$ layer. We are in fact modeling a silicon system with periodic $\delta$ layers every 21.8 nm rather than a single isolated pair of $\delta$ layers. As recently explored by Campbell et al., Campbell et al. (2023) it can be difficult to fully disentangle the impact of $\delta$-doped layers from the high concentration of dopants being simulated. Nonetheless, given the agreement between experimental characterization and DFT calculations of phosphorus $\delta$-doped layers,Miwa et al. (2013); Mazzola et al. (2014a, b); Miwa et al. (2014); Mazzola et al. (2018); Holt et al. (2020) it is reasonable to take these DFT results as good first-order approximations of the electronic structure.
@@ IV Conclusion
Motivated by recent progress in using APAM to create $\delta$-doped layers, we used DFT to predict the electronic structure of both interacting phosphorus and boron $\delta$-doped layers and interacting phosphorus and aluminum $\delta$-doped layers in silicon. At separations of less than $\approx$ 1 nm, we demonstrate that these $\delta$-doped layers will strongly suppress each other, leading to an effective band gap and electronic structure close to that of intrinsic silicon. As the separation between the layers increases to around 1 nm, the suppression decreases, and distinct $\delta$ potentials can be seen at the site of each layer, essentially creating a p-n diode with an intrinsic layer of silicon in the place of the depletion region. This decrease in suppression manifests as a change in the effective band gap that can be tailored by controlling the amount of silicon between the $\delta$-doped layers, ranging from 0.36 eV for 2 nm spacing to 0.86 eV for 0.1 nm spacing. However, the CBM and VBM are spatially separated, so carrier transport between $\delta$-doped layers likely requires tunneling or application of bias. Calculations of the $\delta$-doped layer potentials enable us to predict the tunneling rate between the two layers, finding that the tunneling at higher separations is greater than what would be expected for a pure 1.1 eV triangular barrier. These calculations thus provide a foundation for design of silicon electronics based on interacting $\delta$ layers.
Future work in this area would need to focus on experimentally creating both acceptor and donor $\delta$-doped layers stacked on top of each other in the same silicon device, which has not yet been demonstrated. The growth of clean, relatively defect free, intrinsic silicon between the two $\delta$-doped layers would be a particular challenge. Further work is also needed investigating the electronic structure of $\delta$-doped layers in other group-IV materials such as SiGe, Ge, GeSn, etc. to gain a more complete understanding of the influence of varying strain and host material band structure. Validation of the band structure could be undertaken with angle resolved photoemission spectroscopy (ARPES) as seen in previous experimental work looking at single $\delta$-doped layer silicon structures Miwa et al. (2013, 2014); Strand et al. (2025).
@@ Appendix A Band Structures of isolated single $\delta$-doped layers
The band structure and LDOS of a phosphorus $\delta$-doped layer is shown in Fig. 8, a boron $\delta$-doped layer band structure is shown in Fig. 9, and an aluminum $\delta$-doped layer band structure is shown in Fig. 10. For each structure, the effects are limited entirely to the conduction band (in the case of phosphorus) or the valence band (in the case of boron). Naively, we might expect that the effects of including both a phosphorus and boron $\delta$-doped layer would be a superposition of the phosphorus conduction band and the boron conduction bands. This description only resembles the band structures when the $\delta$-doped layers are separated by $\geq$ 2.0 nm, indicating that interaction between the layers is indeed causing significant suppression of the bands at lower separations. It is notable that SCAN pseudopotentials predict that the $\delta$ potential induced by the phosphorus $\delta$-doped layer is much more significant in magnitude than previous predictions. While the exact magnitude is likely an overestimate, it does indicate that phosphorus $\delta$-doped layers should have an extremely large profile within our calculations. This indicates that the suppression of the phosphorus $\delta$-potential at low separation distances is even more significant than it may appear at first glance.
@@
Figure 11: The doping potential of structures with (a) 0.1 nm, (b) 0.4 nm, (c) 1.0 nm separation between the boron and phosphorus $\delta$-doped layers. The left column displays the potential before the atomic structures are allowed to relax to respond to stress, and the right column is the potential after the structures are allowed to relax.
Figure 12: The doping potential of structures with (a) 2.0 nm, (b) 10.0 nm separation between the boron and phosphorus $\delta$-doped layers. The left column displays the potential before the atomic structures are allowed to relax to respond to stress, and the right column is the potential after the structures are allowed to relax.
To further elucidate the nature of the $\delta$-doped layers interaction, we plot the doping potentials for each separation distance in Figs. 11 and  12. These potentials are calculated by subtracting the planar averaged electrostatic potential of a pure silicon supercell from the planar averaged electrostatic potential of the system with boron and phosphorus $\delta$-doped layers. The doping potentials calculated in the left column of each figure represent the potential before the system is allowed to relax in response to the stress that the doping atoms induce in the system. These graphs more or less follow the pattern that would be expected from the LDOS calculated in Figs. 2 and 3. In the right column, however, we display the potential after the system is allowed to relax in response to stress induced by the doping atoms. These potentials have a much more oscillatory character around the $\delta$-doped layers, which is a direct result of atomic movement in response to the strain slightly altering the periodicity of the nearby silicon atoms. Similar to what is seen in Ref. Campbell et al., 2023, boron $\delta$-doped layers feature particularly strong oscillation with the resulting $\delta$ potential being less well defined within the material compared to the phosphorus $\delta$-doped layers. This is especially clear in the well separated structures at 2.0 and 10.0 nm separation shown in Fig. 12.
Figure 13: The doping potential of Al-P structures with (a) 0.1 nm, (b) 0.4 nm, and (c) 1.0 nm separation between the boron and phosphorus $\delta$-doped layers. The left column displays the potential before the atomic structures are allowed to relax to respond to stress, and the right column is the potential after the structures are allowed to relax.
Figure 14: The doping potential of Al-P structures with (a) 2.0 nm, (b) 10.0 nm separation between the boron and phosphorus $\delta$-doped layers. The left column displays the potential before the atomic structures are allowed to relax to respond to stress, and the right column is the potential after the structures are allowed to relax.
@@
To provide a quantitative comparison between the two systems, we define a decay width $\sigma_{doping}$,
$\sigma_{doping}=z_{thr}-z_{B/Al},$ (3)
where $z_{B/Al}$ is the z position where the B or Al $\delta$ layer is centered, and $z_{thr}$ is the threshold z position, at which the potential oscillation is << than a specified value. We choose the threshold to be a z value with potential oscillations << 0.02 V. We find that for B $\delta$ layers the average $\sigma_{doping}=2.2$ nm, whereas for Al $\delta$ layers, the average $\sigma_{doping}=1.6$ nm. This indicates that B $\delta$ layers induce atomic displacement in the surrounding silicon an average of 0.6 nm further away than Al layers do, which may serve as a partial explanation for why the Al-P structures tend to have more defined peaks in the LDOS than B-P structures do.
Appendix C Tunneling potentials for the boron and aluminum $\delta$-doped layers interacting with a phosphorus $\delta$-doped layers
From the LDOS diagrams in Figures 2-5, we can extract the outline of the conduction band minima (and corresponding valence band maxima with a potential shift) looks like. We then use this as an input potential for the tunneling calculations described in Sec. IIIC. We show the extracted $V(z)$ plots for each system used in our tunneling calculations in Fig. 15.
Figure 15: The conduction band potential of B-P and Al-P structures with (a) 0.1 nm, (b) 0.4 nm, (c) 1.0 nm, (d) 2.0 nm separation between the boron and phosphorus $\delta$-doped layers. These calculations are all for relaxed structures.